# Quantifying Edits Decay in Fine-tuned LLMs

### v1

@@ Abstract
Knowledge editing has emerged as a lightweight alternative to retraining for correcting or injecting specific facts in large language models (LLMs). Meanwhile, fine-tuning remains the default operation for adapting LLMs to new domains and tasks. Despite their widespread adoption, these two post-training interventions have been studied in isolation, leaving open a crucial question: if we fine-tune an edited model, do the edits survive? This question is motivated by two practical scenarios: removing covert or malicious edits, and preserving beneficial edits. If fine-tuning impairs edits as shown in Figure 1, current KE methods become less useful, as every fine-tuned model would require re-editing, which significantly increases the cost; if edits persist, fine-tuned models risk propagating hidden malicious edits, raising serious safety concerns. To this end, we systematically quantify edits decay after fine-tuning, investigating how fine-tuning affects knowledge editing. We evaluate two state-of-the-art editing methods (MEMIT, AlphaEdit) and three fine-tuning approaches (full-parameter, LoRA, DoRA) across five LLMs and three datasets, yielding 232 experimental configurations. Our results show that edits decay after fine-tuning, with survival varying across configurations, e.g., AlphaEdit edits decay more than MEMIT edits. Further, we propose selective-layer fine-tuning and find that fine-tuning edited layers only can effectively remove edits, though at a slight cost to downstream performance. Surprisingly, fine-tuning non-edited layers impairs more edits than full fine-tuning. Overall, our study establishes empirical baselines and actionable strategies for integrating knowledge editing with fine-tuning, and underscores that evaluating model editing requires considering the full LLM application pipeline.
@@ 1 Introduction
Large Language Models (LLMs) can be updated after pre-training through two main approaches. The first is fine-tuning (FT), where model parameters are updated by training the model on a task-specific dataset (Howard & Ruder, 2018). FT also includes parameter-efficient variants such as LoRA (Hu et al., 2022) and DoRA (Liu et al., 2024), collectively referred to as Parameter-Efficient Fine-Tuning (PEFT). The second approach is knowledge editing (KE) (Mazzia et al., 2025). Unlike FT, which adapts a model to specific tasks, KE is used to update the model’s factual knowledge with limited data and compute budget.
Figure 1: An illustration of an LLM ($M$) that undergoes an edit ($M_{ed}$ ) and then fine-tuning ($M_{ed\_ft}$). This process results in the loss of edited knowledge and the production of incorrect outputs. Here is an illustrative example, not real case we found during experiments.
@@ 2.2 Fine-Tuning and Parameter-Efficient Adaptation
FT is the default route for adapting foundation models to domains and tasks (e.g., ULMFiT; Howard & Ruder, 2018). PEFT techniques have become the practical workhorse across the LLM production pipeline: LoRA injects low-rank adapters into frozen backbones (Hu et al., 2022), DoRA decomposes weights into magnitude and direction to better match full-FT capacity (Liu et al., 2024), and adapter families provide modular, swappable components for rapid specialization (Hu et al., 2023). Empirically, PEFT often outperforms few-shot ICL while being dramatically cheaper than full FT (Liu et al., 2022). As a result, PEFT has become the de facto standard across the production pipeline of LLM-based applications. In practice, cloud providers (e.g., Azure [^1] https://learn.microsoft.com/en-us/azure/ai-foundry/concepts/fine-tuning-overview and Google Cloud [^2] https://cloud.google.com/vertex-ai/generative-ai/docs/models/tune-models) and model hubs (e.g. HuggingFace) distribute LoRA/adapter checkpoints or supporting pipeline as compact add-ons, enabling organizations to maintain a single backbone and compose task- or client-specific adapters at deploy time (Hu et al., 2022; Ye et al., 2023; Liu et al., 2024). This widespread industrial adoption makes understanding PEFT’s interaction with KE highly consequential.
@@ 3 Experiment
As discussed above, we are motivated to understand the impact of FT on model editing. To this end, we construct four groups of models: ① base models, $M$, no FT or KE. ② FT-only models, $M_{\text{ft}}$, (we further use $M_{\text{full}}$, $M_{\text{LoRA}}$ and $M_{\text{DoRA}}$ to refer to full size FT, and FT with LoRA and DoRA, respectively); ③ KE-only models, $M_{\text{ed}}$; ④ KE-then-FT models, $M_{\text{ed\_ft}}$. We compare the editing performance gap between $M_{\text{ed}}$ and $M_{\text{ed\_ft}}$ to assess the impact of FT on KE (Section 4.1). We evaluate the downstream task performance of $M_{\text{ft}}$ to validate the FT performance in general (Section 4.5). Our experiments cover five models, two KE datasets, two KE methods, four editing number settings, and four fine-tuning settings, resulting in 216 independent model configurations. Following Liu et al. (2024), we use the Commonsense dataset as the FT corpus and evaluate the model on eight downstream tasks (Section 3.1). We summarise all configurations in Table 9 in the Appendix B.
@@ 3.1 Models and datasets
Following previous works on KE (Meng et al., 2022; Fang et al., 2025), we include GPT-J-6B (denoted as GPT-JWang & Komatsuzaki (2021)), GPT2-XL (Radford et al., 2019)), Llama2-7B-hf (denoted as Llama2Touvron et al. (2023)), and Llama3.1-8B-Instruct (denoted as Llama3Grattafiori et al. (2024)). We initially considered using DeepSeek as it is a recent SOTA model. However, its poor KE performance (see Section 4.4) renders the analysis of further fine-tuning uninformative. Details of models can be referred to in Table 8.
@@
Following Meng et al. (2022)’s work, we use COUNTERFACT (Meng et al., 2022) and zsRE (Levy et al., 2017; Mitchell et al., 2022b). COUNTERFACT comprises 21,89021{,}890 pairs of factual and counterfactual statements, each paired with multiple paraphrased prompts. zsRE is a question-answering dataset composed of real-world statements. For our experiments, we use its evaluation subset, which comprises 19,08619{,}086 instances, each consisting of a factual statement and an associated paraphrased prompt.
@@
Following Liu et al. (2024), we include the commonsense reasoning dataset (Hu et al. (2023)) for fine-tuning. This dataset comprises 170,000 data points, consisting of eight downstream tasks: BoolQ, PIQA, SIQA, HellaSwag, WinoGrande, ARC-e, ARC-c, and OBQA. All downstream tasks are multiple-choice questions, except for BoolQ, which contains yes-or-no questions.
@@
We focus on two popular parameter-modifying methods: MEMIT (Meng et al., 2023) and AlphaEdit (Fang et al., 2025). MEMIT enables multiple knowledge insertions simultaneously by employing matrix optimisation. AlphaEdit projects perturbation into the null space to ensure it does not affect other facts. We follow the editing hyper-parameter settings as reported in Meng et al. (2023); Fang et al. (2025). [^3] Our reproduction of MEMIT and AlphaEdit yields results that differ slightly from those reported in their original papers, but the differences fall within the reported standard deviations. Detailed results are provided in Appendix C.
@@
We evaluate KE performance using three metrics: Efficacy Success (ES), Paraphrase Success (PS) and Neighborhood Success, (NS) which measure KE success rate with the provided prompts, success rate with paraphrased prompts, and the level of influence on irrelevant facts, respectively (particularized in Appendix D). They all range from 0 to 1, with 1 representing the best and 0 the worst.
Moreover, we introduce Edit Flip Ratio (EFR) to quantify, at the individual case level, how many successful edits (success on ES) become unsuccessful after fine-tuning. Unlike ES, which measures the overall proportion of successful edits on the full test set after editing, i.e., the denominator is the number of total target edits, EFR conditions on successful edits and computes the fraction that are undone after fine-tuning, i.e., the denominator is the number of successful edits. EFR directly addresses our research question by quantifying the persistence of edits under fine-tuning.
As demonstrated in Formula 1, $s_{i}$ is a binary flag recording an edit is successfully edited, where $\mathbb{P}$ represents the probability of the model predicting the editing target. We then define a flip case (note as $i$) as an edit that is successful after editing ($s_{i}^{\text{ed}}$ = 1), but becomes unsuccessful after fine-tuning ($s_{i}^{\text{ed\_ft}}$ = 0). EFR is the proportion of flip cases within total edits $\mathcal{B}$ as shown in Formula 1. A higher EFR indicates that fine-tuning impairs more successful edits.
$\text{EFR}=\frac{\left|\left\{i\in\mathcal{B}\ \middle|\ s_{i}^{\text{ed}}=1\ \wedge\ s_{i}^{\text{ed\_ft}}=0\right\}\right|}{|\mathcal{B}|}\quad s_{i}=\begin{cases}1&\text{if }\mathbb{P}(\text{editing target})>\mathbb{P}(\text{original target})\\ 0&\text{otherwise}\end{cases}$ (1)
For fine-tuning, we evaluate the model on the eight downstream tasks described in Section 3.1, which test its reasoning abilities across diverse domains such as physics, social implications, and scientific knowledge (Hu et al., 2023).
@@ 4.1 Editing Performance after Fine-tuning
Table 1 presents the editing success rate (ES) of GPT-J, GPT2-XL, and Llama2 on zsRE, before and after fine-tuning. Results for the remaining metrics, i.e., Paraphrase Success and Specificity Success, as well as the results on the COUNTERFACT dataset, are provided in Table 11, Table 12, and Table 13 in the Appendix E. We also present the performance of editing-only ($M_{\text{ed}}$) and fine-tuning-only ($M_{\text{ft}}$) in the Table 18-Table 29, validating the setup.
Table 1: Success rate (ES, %) of edited GPT-J, GPT2-XL and Llama2 between before and after fine-tuning. Full results across KE metrics are in Tables 11-13.
@@
Overall, fine-tuning reduces editing performance, with only four exceptions showing comparable editing performance to $M_{\text{ed}}$ (“No ft” in Table 4.1). For example, when applying MEMIT to edit 1000 facts on GPT2-XL, after full fine-tuning, the ES increases by $3.37$ percentage points (p.p.). This is counterintuitive, as some target edits that initially failed became successful after fine-tuning. We further analyze this in Section 4.2. Among all decay cases, LoRA fine-tuning on Llama2 after 1000 edits on zsRE using AlphaEdit brings the largest decrease in editing performance, from $93.23\%$ to $50.45\%$. Additionally, while fine-tuning impairs KE performance, the extent of this effect varies across KE setups, fine-tuning configurations, and models, as discussed below. We do not conduct fine-tuning on DeepSeek, as its KE performance is substantially lower than that of other models, rendering it impractical for further adoption in downstream tasks. Additional discussion is provided in Section 4.4.
@@
We find that full fine-tuning impairs a markedly larger fraction of edits than LoRA and DoRA, as shown by the average decrease across all models in Table 2: $38.10\%$ for full fine-tuning, versus $28.71\%$ and $29.88\%$ for LoRA and DoRA. DoRA demonstrates a slightly stronger ability to remove edits than LoRA. This pattern varies with model architecture and edit scale. As shown in Table 1, Llama2’s edit success rate decreases sharply from $86.03\%$ to $22.67\%$ under full fine-tuning, whereas LoRA and DoRA yield considerably smaller declines of $9.73$ and $14.03$ p.p., respectively.
@@
Between AlphaEdit and MEMIT, AlphaEdit exhibits greater decay after FT, i.e., its edits are more easily removed. Take Llama2 as an example, LoRA fine-tuning reduces MEMIT performance by $9.73$ p.p., compared to $35.37$ p.p. for AlphaEdit. When the number of edits increases to 10,00010{,}000, the performance gap widens to $33.50$ p.p., indicating that large-scale edits exacerbate AlphaEdit’s vulnerability. Such a pattern is observed consistently across GPT-J and GPT2-XL.
@@
As shown by the  Overall Avg. in Table 2, GPT-J is the most stable under fine-tuning, followed by Llama2, whereas GPT2-XL exhibits the largest variability. GPT-J achieves the smallest average decrease ($15.93\%$) compared to Llama2 ($39.92\%$) and GPT2-XL ($40.84\%$), and also has the lowest standard deviation ($19.04\%$), indicating more consistent degradation across KE and fine-tuning methods. Although GPT2-XL records the highest average decrease ($40.84\%$), its standard deviation is slightly lower than Llama2’s, suggesting marginally greater stability under fine-tuning.
@@
As illustrated in Figure 2, models edited with zsRE generally experience a greater decline in KE performance compared to those edited with COUNTERFACT, regardless of the number of edits or the KE method employed. For example, in the third subplot, where both models are edited using MEMIT with 10,00010{,}000 edits, the ES of the zsRE-edited model decreases by $8$ p.p., whereas the COUNTERFACT-edited model experiences only a $2$ p.p. decrease. The larger performance drop on zsRE likely stems from differences in evaluation metrics: COUNTERFACT deems an edit successful if the edited object has a higher probability than the original, whereas zsRE requires the model to generate the edited object correctly, a stricter criterion.
Edit Flip Ratio

### v3

# Can Fine-Tuning Erase Your Edits? On the Fragile Coexistence of Knowledge Editing and Adaptation

@@ Abstract
Knowledge editing has emerged as a lightweight alternative to retraining for correcting or injecting specific facts in large language models (LLMs). Meanwhile, fine-tuning remains the default operation for adapting LLMs to new domains and tasks. Despite their widespread adoption, these two post-training interventions have been studied in isolation, leaving open a crucial question: if we fine-tune an edited model, do the edits survive? This question is motivated by two practical scenarios: removing covert or malicious edits, and preserving beneficial edits. If fine-tuning impairs edits (Fig. 1), current KE methods become less useful, as every fine-tuned model would require re-editing, which significantly increases the cost; if edits persist, fine-tuned models risk propagating hidden malicious edits, raising serious safety concerns. To this end, we systematically quantify edit decay after fine-tuning, investigating how fine-tuning affects knowledge editing. Our results show that edits decay after fine-tuning, with survival varying across configurations, e.g., AlphaEdit edits decay more than MEMIT edits. Further, we find that fine-tuning edited layers only can effectively remove edits, though at a slight cost to downstream performance. Surprisingly, fine-tuning non-edited layers impairs more edits than full fine-tuning. Overall, our study establishes empirical baselines and actionable strategies for integrating knowledge editing with fine-tuning, and underscores that evaluating model editing requires considering the full LLM application pipeline.
Can Fine-Tuning Erase Your Edits? On the Fragile Coexistence of Knowledge Editing and Adaptation
Yinjie Cheng1, Paul Youssef2, Christin Seifert2, Jörg Schlötterer2, Zhixue Zhao1 1University of Sheffield  2Phillips-Universität Marburg {ycheng80, zhixue.zhao}@sheffield.ac.uk {paul.youssef, christin.seifert, joerg.schloetterer}@uni-marburg.de
@@ 1 Introduction
Large Language Models (LLMs) can be updated after pre-training through two main approaches. The first is fine-tuning (FT), where model parameters are updated by training the model on a task-specific dataset (Howard and Ruder, 2018). FT also includes parameter-efficient variants such as LoRA (Hu et al., 2022) and DoRA (Liu et al., 2024), collectively referred to as Parameter-Efficient Fine-Tuning (PEFT). The second approach is knowledge editing (KE) (Mazzia et al., 2025). Unlike FT, which adapts a model to specific tasks, KE is used to update the model’s factual knowledge with limited data and compute budget.
@@
Figure 1: An illustration of an LLM ($M$) that undergoes an edit ($M_{ed}$ ) and then fine-tuning ($M_{ed\_ft}$). This process results in the loss of edited knowledge and the production of incorrect outputs. Here is an illustrative example, we show real cases in Sec. 4.2.
@@ 2.2 Fine-Tuning and Parameter-Efficient Adaptation
FT is the default route for adapting foundation models to domains and tasks (e.g., ULMFiT; Howard and Ruder, 2018). PEFT techniques have become the practical workhorse across the LLM production pipeline: LoRA injects low-rank adapters into frozen backbones (Hu et al., 2022), DoRA decomposes weights into magnitude and direction to better match full-FT capacity (Liu et al., 2024), and adapter families provide modular, swappable components for rapid specialization (Hu et al., 2023). Empirically, PEFT often outperforms few-shot ICL while being dramatically cheaper than full FT (Liu et al., 2022). As a result, PEFT has become the de facto standard across the production pipeline of LLM-based applications. In practice, cloud providers (e.g., Azure [^1] https://learn.microsoft.com/en-us/azure/ai-foundry/concepts/fine-tuning-overview and Google Cloud [^2] https://cloud.google.com/vertex-ai/generative-ai/docs/models/tune-models) and model hubs (e.g. HuggingFace) distribute LoRA/adapter checkpoints or supporting pipeline as compact add-ons, enabling organizations to maintain a single backbone and compose task- or client-specific adapters at deploy time (Hu et al., 2022; Ye et al., 2023; Liu et al., 2024). This widespread industrial adoption makes understanding PEFT’s interaction with KE highly consequential.
@@ 3 Experiment
As discussed above, we are motivated to understand the impact of FT on model editing. To this end, we construct four groups of models: ① base models, $M$, no FT or KE. ② FT-only models, $M_{\text{ft}}$, (we further use $M_{\text{full}}$, $M_{\text{LoRA}}$ and $M_{\text{DoRA}}$ to refer to full size FT, and FT with LoRA and DoRA, respectively); ③ KE-only models, $M_{\text{ed}}$; ④ KE-then-FT models, $M_{\text{ed\_ft}}$. We compare the editing performance gap between $M_{\text{ed}}$ and $M_{\text{ed\_ft}}$ to assess the impact of FT on KE (Sec. 4.1). We evaluate the downstream task performance of $M_{\text{ft}}$ to validate the FT performance in general (Sec. 4.5). Our experiments cover five models, two KE datasets, two KE methods, four editing number settings, and four fine-tuning settings, resulting in 216 independent model configurations. Following Liu et al. (2024), we use the Commonsense dataset as the FT corpus and evaluate the model on eight downstream tasks (Sec. 3.1). We summarise all configurations in Tab. 8 in the App. B.
@@ 3.1 Models and datasets
Following previous works on KE (Meng et al., 2022; Fang et al., 2025), we include GPT-J-6B (denoted as GPT-JWang and Komatsuzaki (2021)), GPT2-XL (Radford et al., 2019)), Llama2-7B-hf (denoted as Llama2Touvron et al. (2023)), and Llama3.1-8B-Instruct (denoted as Llama3Grattafiori et al. (2024)). We initially considered using DeepSeek as it is a recent SOTA model. However, its poor KE performance (see Sec. 4.4) renders the analysis of further fine-tuning uninformative. Details of models can be referred to in Tab. 7.
@@
Following Meng et al. (2022)’s work, we use COUNTERFACT (Meng et al., 2022) and zsRE (Levy et al., 2017; Mitchell et al., 2022b). COUNTERFACT comprises 21,91921{,}919 pairs of factual and counterfactual statements, each paired with multiple paraphrased prompts. zsRE is a question-answering dataset drawn from real-world knowledge sources like Wikipedia and Wikidata. For our experiments, we use its evaluation subset, which comprises $19,086$ instances, each consisting of a factual statement and an associated paraphrased prompt.
@@
Following Liu et al. (2024), we include the commonsense reasoning dataset (Hu et al. (2023)) for fine-tuning. This dataset comprises $170,000$ data points, consisting of eight downstream tasks: BoolQ, PIQA, SIQA, HellaSwag, WinoGrande, ARC-e, ARC-c, and OBQA. All downstream tasks are multiple-choice questions, except for BoolQ, which contains yes-or-no questions. We also include HotpotQA, a multi-hop question-answering dataset containing $97,900$ entities, as our fine-tuning dataset. It is used to evaluate the model’s performance when fine-tuned on a knowledge-intensive dataset.
@@
We focus on three popular parameter-modifying methods: MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2025) and MEND (Mitchell et al., 2022a). MEMIT enables multiple knowledge insertions simultaneously by employing matrix optimisation. AlphaEdit projects perturbation into the null space to ensure it does not affect other facts. MEND decomposes the fine-tuning gradient into low-rank patches and applies them as targeted updates to the model’s weights. We follow the editing hyper-parameter settings as reported in Meng et al. (2023); Fang et al. (2025); Mitchell et al. (2022a). [^3] Our reproduction of MEMIT and AlphaEdit yields results that differ slightly from those reported in their original papers, but the differences fall within the reported standard deviations. Detailed results are provided in App. C.
@@
We evaluate KE performance using three metrics: Efficacy Success (ES), Paraphrase Success (PS) and Neighborhood Success, (NS) which measure KE success rate with the provided prompts, success rate with paraphrased prompts, and the level of influence on irrelevant facts, respectively (particularized in App. D). They all range from 0 to 1, with 1 representing the best and 0 the worst.
Moreover, we introduce Edit Flip Ratio (EFR) as a metric to quantify, at the individual edit level, the number of successful edits that become unsuccessful after fine-tuning. Specifically, EFR exclusively measures the stability of succeeded edits under fine-tuning and addresses a key limitation of the ES metric, which measures only overall knowledge-editing performance.
We use a binary indicator $s^{\text{M}}_{i}$ to represent the editing outcome for fact $i$ in model $M$. Indicator $s^{\text{M}}_{i}$ takes values in $\{0,1\}$, where 1 indicates that the edit is successful, and 0 otherwise [^4] For instance, $s^{\text{ed\_ft}}_{i}=1$ signifies the $i^{\text{th}}$ edit remains successful after fine-tuning in model $M_{\text{ed\_ft}}$.. The evaluation criteria for the indicator are consistent with KE metrics (Equ. 7-8 in App. D). We then define a flipped case as an edit that is successful after KE ($s^{\mathrm{M_{\text{ed}}}}_{i}=1$), but becomes unsuccessful after fine-tuning ($s^{\mathrm{M_{\text{ed\_ft}}}}_{i}=0$). Accordingly, EFR is the probability of having flip cases in $M_{\text{ed\_ft}}$, as shown in Equ. 1
$\mathrm{EFR}=\Pr\left(s_{i}^{\mathrm{M_{\text{ed\_ft}}}}=0\,\middle|\,s_{i}^{\mathrm{M_{\text{ed}}}}=1\right)$ (1)
For fine-tuning, we evaluate the model on the eight downstream tasks described in Sec. 3.1, which test its reasoning abilities across diverse domains such as physics, social implications, and scientific knowledge (Hu et al., 2023).
@@ 4.1 Editing Performance after Fine-tuning
Tab. 1 presents the editing success rate (ES) of GPT-J, GPT2-XL, and Llama2 on zsRE, before and after fine-tuning on Commonsense datasets. Results for the remaining metrics (PS and NS), as well as the results on other datasets (COUNTERFACT and HotpotQA) or using MEND, are provided in Tab. 10 to 15 in Appendix E. We also present the performance of editing-only ($M_{\text{ed}}$) and fine-tuning-only ($M_{\text{ft}}$) in the Tab. 21-Tab. 34, validating the setup.
@@
Table 1: Success rate (ES, %) of edited GPT-J, GPT2-XL and Llama2 between before and after fine-tuning. Full results across KE metrics are in Tables 10-12.
Overall, fine-tuning reduces editing performance, with only four exceptions showing comparable editing performance to $M_{\text{ed}}$ (“No ft” in Tab. 4.1). For example, when applying MEMIT to edit 1000 facts on GPT2-XL, after full fine-tuning, the ES increases by $3.37$ percentage points (p.p.). This is counterintuitive, as some target edits that initially failed became successful after fine-tuning. We further analyze this in Sec. 4.2. Among all decay cases, LoRA fine-tuning on Llama2 after 1000 edits on zsRE using AlphaEdit brings the largest decrease in editing performance, from $93.23\%$ to $50.45\%$. Additionally, while fine-tuning impairs KE performance, the extent of this effect varies across KE setups, fine-tuning configurations, and models, as discussed below. We do not conduct fine-tuning on DeepSeek, as its KE performance is substantially lower than that of other models, rendering it impractical for further adoption in downstream tasks. Additional discussion is provided in Sec. 4.4.
@@
We find that full fine-tuning impairs a markedly larger fraction of edits than LoRA and DoRA, as shown by the average decrease across all models in Tab. 2: $38.10\%$ for full fine-tuning, versus $28.71\%$ and $29.88\%$ for LoRA and DoRA. DoRA demonstrates a slightly stronger ability to remove edits than LoRA. This pattern varies with model architecture and edit scale. As shown in Tab. 1, Llama2’s edit success rate decreases sharply from $86.03\%$ to $22.67\%$ under full fine-tuning, whereas LoRA and DoRA yield considerably smaller declines of $9.73$ and $14.03$ p.p., respectively.
@@
Between AlphaEdit and MEMIT, AlphaEdit exhibits greater decay after FT, i.e., its edits are more easily removed. Take Llama2 as an example, LoRA fine-tuning reduces MEMIT performance by $9.73$ p.p., compared to $35.37$ p.p. for AlphaEdit. When the number of edits increases to 10,00010{,}000, the performance gap widens to $33.50$ p.p., indicating that large-scale edits exacerbate AlphaEdit’s vulnerability. Such a pattern is observed consistently across GPT-J and GPT2-XL. This phenomenon may stem from the Null-Space Vulnerability of AlphaEdit. By constraining $\Delta W^{M_{\text{ed}}}$ to the null space of Fisher directions (Fang et al., 2025), AlphaEdit reduces interference with existing knowledge but places edits in regions that fine-tuning does not prioritize. Since fine-tuning gradients concentrate along high-curvature directions (Wu et al., 2024), updates orthogonal to these (i.e., in the null space) are unstable and susceptible to shrinkage or rotation, explaining AlphaEdit’s fragility relative to MEMIT. Discussions about MEND is in App. E.
@@
As shown by the  Overall Avg. in Tab. 2, GPT-J is the most stable under fine-tuning, followed by Llama2, whereas GPT2-XL exhibits the largest variability. GPT-J achieves the smallest average decrease ($15.93\%$) compared to Llama2 ($39.92\%$) and GPT2-XL ($40.84\%$), and also has the lowest standard deviation ($19.04\%$), indicating more consistent degradation across KE and fine-tuning methods. Although GPT2-XL records the highest average decrease ($40.84\%$), its standard deviation is slightly lower than Llama2’s, suggesting marginally greater stability under fine-tuning. Besides, Llama3.1’s performance is similar to Llama2, which can be checked in App. H.2
@@
As illustrated in Fig. 2, models edited with zsRE generally experience a greater decline in KE performance compared to those edited with COUNTERFACT, regardless of the number of edits or the KE method employed. For example, in the third subplot, where both models are edited using MEMIT with 10,00010{,}000 edits, the ES of the zsRE-edited model decreases by $8$ p.p., whereas the COUNTERFACT-edited model experiences only a $2$ p.p. decrease. The larger performance drop on zsRE likely stems from differences in evaluation metrics: COUNTERFACT deems an edit successful if the edited object has a higher probability than the original, whereas zsRE requires the model to generate the edited object correctly, a stricter criterion.
Edit Flip Ratio